# Version dated: November 6, 2025

### v1

@@ 1 Introduction
Continuous-time Markov chains (CTMCs) are a probabilistic modeling tool used in a wide range of fields such as phylogenetics, epidemiology, survival analysis, and biochemistry (jukes_evolution_1969; kalbfleisch_analysis_1985). A CTMC is a stochastic process that randomly transitions between discrete states over continuous time. The transitions occur at exponentially distributed waiting times, and the future state depends only on the present state, not on the past (Markov property). The parameters describing the CTMC can be collected into a matrix called an infinitesimal rate matrix or generator. A key inferential task to understand the CTMC dynamics is, therefore, estimating the entries (rates) of such a matrix. This is challenging since the number of rates increases quadratically with the number of discrete states and the data might not carry information on transitions between all pairs of states.
@@
One approach to reducing the number of free parameters in the rate matrix is to impose structural restrictions. This method has been particularly effective in fields such as evolutionary biology, where a lower-dimensional parametrization can be motivated by knowledge of the underlying biological dynamics. A rich body of literature has explored DNA substitution processes, proposing and testing various constraints (jukes_evolution_1969; kimura_simple_1980).
@@
Another modeling technique leverages pre-existing information captured by covariates. One upside of this approach is that, as a byproduct, it sheds light on whether and how those covariates influence the CTMC jumping process. For instance, phylogeography explores the emergence and dispersal of biological entities, including pathogenic viruses, using CTMCs (pybus_unifying_2012; lemey_unifying_2014). The rate of a viral infection spreading from one country to another is likely dependent on the distance that separates the two countries as well as their population sizes.
Within a frequentist framework, kalbfleisch_analysis_1985 proposed to study the relationship between covariates and rates based on a log-linear model. This approach was then extended to a Bayesian setting in the context of phylogeography. Motivated by a rich set of examples, researchers have actively contributed to the improvement of these models. Nonetheless, the geographic locations under study tend to be high in number, and finding a credible direct parametrization that describes the movement among them remains challenging.
@@
There has been a flourishing literature on Markov chain Monte Carlo (MCMC) techniques (metropolis_equation_1953; hastings_monte_1970) to sample from the posterior distribution of CTMC rates (suchard_bayesian_2001). However, when the state space is large, employing a simple Metropolis-Hastings-based MCMC for sampling rates becomes challenging, not only due to their high dimensionality but also because their strong mutual dependencies hinder efficient exploration of the parameter space. Adopting a Hamiltonian Monte Carlo (HMC)-based sampling scheme (neal_mcmc_2011) has proven to be an effective solution for sampling high-dimensional and highly dependent parameters. For instance, in the context of CTMCs, zhao_bayesian_2016 use HMC to sample rates based on a Bayesian log-linear model.
@@
A major challenge of HMC-based sampling for many CTMC models is the repeated computation of the (log-) likelihood gradient with respect to the parameters of interest. For partially observed CTMCs, this involves evaluating multiple matrix exponential derivatives (computational complexity ${\cal O}(K^{3})$) for each of the infinitesimal rates ($O(K^{2})$), leading to a total complexity of $O(K^{5})$ operations and creating a severe computational bottleneck even for moderately-sized state spaces. To address this, we explore an approximation of the matrix exponential gradient that reduces the overall cost to only $O(K^{2})$ (didier_surprising_2024), gaining a three orders-of-magnitude improvement. Crucially, using an approximate gradient does not hinder HMC convergence, as supported by the literature on HMC using surrogate trajectories (li_neural_2019). Finally, we adapt the proposed methods to Bayesian phylogeographic inference following magee_random-effects_2024, and demonstrate how the aforementioned approximation is even more critical for efficient sampling under flexible GP modeling assumptions.
@@ 2.1 Brief introduction to continuous-time Markov chains
More intuitively, if the CTMC $Y_{t}$ starts at some state $i$, it will leave such state after an exponentially distributed time with rate $q_{i}$ (time component) and will jump to a different state $j$ with probability $\frac{q_{ij}}{q_{i}}$ (location component).
The exponential random variable describing the time spent in a state $i$ before leaving it has been assigned several names; in this work, we refer to it as the waiting time, $W_{i}$. In addition, we denote the realized value of the CTMC $Y(t)$ at time $t$ with $y(t)$. If we assume the CTMC is observed $N$ times, then we refer to the vector of observations as $\mathbf{y}=(y_{1},y_{2},\dots,y_{N})$. Similarly, we refer to the times when each value $y_{i}$ is observed as $t_{i}$; we also collect those values in the vector $\mathbf{t}=(t_{1},t_{2},\dots,t_{N})$.
@@ 2.2.1 Fully observed data
${p}\hskip-1.00006pt\left(\mathbf{y},\mathbf{t}\hskip-1.99997pt\mid\hskip-1.00006pt\mathbf{Q},\text{fully observed}\right)=\pi^{*}(y_{1})\prod_{i=2}^{N}\left[{\mathbb{P}}\hskip-1.99997pt\left(W_{y_{i}}=t_{i}-t_{i-1}\right)\frac{q_{y_{i-1},y_{i}}}{q_{y_{i}}}\right],$ (2)
@@ 2.2.2 Partially observed data
A more realistic scenario involves a partially observed CTMC. Specifically, the simplest case occurs when the state of the CTMC is known at time 0 and at time $t$, while its behavior throughout the interval $(0,t)$ remains unknown. This implies that an object of major importance is the finite-time transition probability of starting at a time, say 0, on state $i$ and being on state $j$ at a later time, say $t$, that is:
@@
Collecting all such values in a matrix whose rows represent the starting states and columns the ending states, we obtain the finite-time transition probability matrix ${\mathbf{P}}\hskip-1.99997pt\left(t\right)=\left(P_{ij}(t)\right)_{ij}$. If we assume the identity matrix $\mathbf{I}$ as the initial condition of the system, that is the system is in its current state without any transitions at time $t=0$, then it is easy to derive a closed-form formula for ${\mathbf{P}}\hskip-1.99997pt\left(t\right)$ based on the infinitesimal rate matrix $\mathbf{Q}$:
${\mathbf{P}}\hskip-1.99997pt\left(t\right)=e^{t\mathbf{Q}},$ (4)
where the matrix exponential is defined as follows:
$e^{t\mathbf{Q}}:=\sum_{i=0}^{\infty}\frac{(t\mathbf{Q})^{i}}{i!}.$ (5)
Now we consider three cases where the data are only partially observed. We begin by assuming that we observe the data sequentially. Within this framework, we distinguish when the sampling times are known and when they are not. Then, we briefly introduce the more general case where observations are not assumed to be sequential.
@@
The simplest partially observed scenario arises when the sequence of observations includes both the states and the corresponding sampling times $(\mathbf{y},\mathbf{t})$, but there is no guarantee that all the jumps of the CTMC are observed.
Then, we can write the joint distribution of $(\mathbf{y},\mathbf{t})$ using the Markov property and the finite transition probabilities:
${p}\hskip-1.00006pt\left(\mathbf{y},\mathbf{t}\hskip-1.99997pt\mid\hskip-1.00006pt\mathbf{Q}\right)=\pi^{*}(y_{1})\prod_{i=2}^{N}P_{y_{i-1},y_{i}}(t_{i}).$ (6)
With the exception of specific, mostly low-dimensional cases, such as the Jukes-Cantor (JK) substitution model (jukes_evolution_1969) for DNA mutations, the values $P_{y_{i-1},y_{i}}(t_{i})$ are generally not available in closed-form. However, they can be computed by evaluating the matrix exponential for the specified time interval, which requires a computational effort of cubic-order (${\cal O}(K^{3})$).
@@
Suppose instead that the realizations in the vector $\mathbf{y}$ are sequentially observed, but the observation times are unknown. From formula 4 it is clear that the waiting times and the infinitesimal rate matrix are jointly identifiable only up to a multiplicative factor. A candidate solution is to constrain the structure of $\mathbf{Q}$ and define a new time scale dependent on the infinitesimal rate matrix.
To achieve this, let $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{K})$ be a probability mass vector on the $K$ states in $\mathbf{S}$ which represents the a priori state frequencies; $\boldsymbol{\pi}$ can represent $\mathbf{Q}$’s stationary distribution ($\boldsymbol{\pi}=\boldsymbol{\pi}^{*}$), but this need not be the case. Let $N_{\mathbf{Q}}(t)$ count the (random) number of jumps during the period $[0,t]$ with respect to the probability vector $\boldsymbol{\pi}$. It can be shown that its expected value is equal to:
$\mathbb{E}\left[N_{\mathbf{Q}}(t)\right]=t\sum_{i\in\mathbf{S}}\pi_{i}q_{i},$ (7)
recalling that $q_{i}=\sum_{j\neq i}q_{ij}$. A way to ensure this expected value depends only on time is to divide each element of the rate matrix $\mathbf{Q}$ by a normalizing constant $\psi$ equal to:
@@
Therefore, defining the matrix $\mathbf{\Lambda}=(\lambda_{ij})_{ij}$ as the normalized version of $\mathbf{Q}$, that is
$\lambda_{ij}=\frac{1}{\psi}q_{ij}\qquad\text{for every }i,j\in\left\{1,2,\dots,K\right\},$ (9)
we have created a new CTMC where the expected number of jumps during a period $[0,t]$ is equal to the time:
$\mathbb{E}\left[N_{\mathbf{\Lambda}}(t)\right]=t.$ (10)
Using the normalized matrix constrains the rate values to satisfy Condition 10, solving the identifiability problem. Indeed, suppose we respectively multiply and divide the matrix $\mathbf{\Lambda}$ and time by a scale factor, then the scale factor disappears in the normalization step, implying that after the transformation, the normalized rate matrix is unchanged, while the time variable has changed.
This approach suggests that we can, and should, reinterpret time in terms of the expected number of jumps of the CTMC. For example, a unit time length is equal to the time necessary for the normalized CTMC to experience on average one jump ($\mathbb{E}\left[N_{\mathbf{\Lambda}}(1)\right]=1$ by construction). In the phylogenetic literature, a field that heavily relies on CTMCs, this time scale is often referred to as the “evolutionary time” scale, as opposed to the natural time scale. It provides a standardized way to compare processes that may evolve at different speeds.
Assigning a prior ${p}\hskip-1.00006pt\left(\mathbf{t}\right)$ to the times, the likelihood of the observed data conditional on the rate matrix is:
@@
where ${p}\hskip-1.00006pt\left(\mathbf{y},\mathbf{t}\hskip-1.99997pt\mid\hskip-1.00006pt\mathbf{Q}\right)$ can be computed as in Equation 6.
@@ 2.3 Prior on the infinitesimal rate matrix
The non-negativity constraints on the non-diagonal elements of the rate matrix $\mathbf{Q}$, if directly modeled, prevent the use of convenient prior distributions defined on the real line and complicate the use of gradient-based samplers. As is customary, we introduce a vector of auxiliary parameters $\boldsymbol{\theta}=(\theta_{1},\theta_{2},\dots,\theta_{K})$ and link them to the rates by a non-negative differentiable one-to-one transformation $g(\cdot)$, such that:
@@
To relax the restrictive assumptions of the log-linear model employed in the previous literature (pybus_unifying_2012; lemey_unifying_2014; zhao_bayesian_2016), we propose to model the relationship between the parameters and the covariate using a more flexible GP prior. This allows us to capture complex, nonlinear dependencies that a log-linear framework cannot adequately represent.
@@
The GP prior is a powerful Bayesian nonparametric prior that defines a distribution over more generic functions. Beyond accommodating nonlinearities, the GP enables a principled way to model uncertainty, as it provides not only point estimates but also credible intervals for inferred relationships. The GP is parameterized by a mean and a covariance function. The covariate function, also known as the GP kernel, defines the dependencies between function values at different input points, controlling both the smoothness and variability of the inferred function. By specifying an appropriate kernel, we can encode prior beliefs about the structure of the relationship, such as stationarity, periodicity, or varying degrees of smoothness. Common choices include the squared exponential kernel, which enforces smooth variations, and the Matérn kernel, which allows for more flexible roughness.
@@
If we consider a set of $P$ covariates $(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{P})$, then we can model the parameters $\theta_{ij}$ with an additive GP prior (duvenaud_additive_2011), that is:
@@
where all $f_{p}$’s are independent and allowed to have different kernels $\mathbf{k}_{p}(\cdot,\cdot|\boldsymbol{\alpha}_{p})$. This last expression can then be integrated out leveraging the independence of the $f_{p}$’s. Specifically, in vector form, we obtain:
$\boldsymbol{\theta}$ $\sim\mathcal{GP}(\boldsymbol{0},\sum_{p=1}^{P}\mathbf{k}_{p}(\mathbf{x}_{p},\mathbf{x}_{p}^{\prime}|\boldsymbol{\alpha}_{p})).$ (15)
This framework can then be easily generalized to account for interactions between covariates (duvenaud_additive_2011).

### v2

@@ 1 Introduction
Continuous-time Markov chains (CTMCs) are a probabilistic modeling tool used in a wide range of fields such as phylogenetics, epidemiology, survival analysis, and biochemistry (Jukes and Cantor, 1969; Kalbfleisch and Lawless, 1985). A CTMC is a stochastic process that randomly transitions between discrete states over continuous time. The transitions occur at exponentially distributed waiting times, and the future state depends only on the present state, not on the past (Markov property). The parameters describing the CTMC can be collected into a matrix called an infinitesimal rate matrix or generator. A key inferential task to understand the CTMC dynamics is, therefore, estimating the entries (rates) of such a matrix. This is challenging since the number of rates increases quadratically with the number of discrete states and the data might not carry information on transitions between all pairs of states.
@@
One approach to reducing the number of free parameters in the rate matrix is to impose structural restrictions. This method has been particularly effective in fields such as evolutionary biology, where a lower-dimensional parametrization can be motivated by knowledge of the underlying biological dynamics. A rich body of literature has explored DNA substitution processes, proposing and testing various constraints (Jukes and Cantor, 1969; Kimura, 1980).
@@
Another modeling technique leverages pre-existing information captured by covariates. One upside of this approach is that, as a byproduct, it sheds light on whether and how those covariates influence the CTMC jumping process. For instance, phylogeography explores the emergence and dispersal of biological entities, including pathogenic viruses, using CTMCs (Pybus et al., 2012; Lemey et al., 2014). The rate of a viral infection spreading from one country to another is likely dependent on the distance that separates the two countries as well as their population sizes.
Within a frequentist framework, Kalbfleisch and Lawless (1985) proposed to study the relationship between covariates and rates based on a log-linear model. This approach was then extended to a Bayesian setting in the context of phylogeography. Motivated by a rich set of examples, researchers have actively contributed to the improvement of these models. Nonetheless, the geographic locations under study tend to be high in number, and finding a credible direct parametrization that describes the movement among them remains challenging.
@@
There has been a flourishing literature on Markov chain Monte Carlo (MCMC) techniques (Metropolis et al., 1953; Hastings, 1970) to sample from the posterior distribution of CTMC rates (Suchard et al., 2001). However, when the state space is large, employing a simple Metropolis-Hastings-based MCMC for sampling rates becomes challenging, not only due to their high dimensionality but also because their strong mutual dependencies hinder efficient exploration of the parameter space. Adopting a Hamiltonian Monte Carlo (HMC)-based sampling scheme (Neal, 2011) has proven to be an effective solution for sampling high-dimensional and highly dependent parameters. For instance, in the context of CTMCs, Zhao et al. (2016) use HMC to sample rates based on a Bayesian log-linear model.
@@
A major challenge of HMC-based sampling for many CTMC models is the repeated computation of the (log-) likelihood gradient with respect to the parameters of interest. For partially observed CTMCs, this involves evaluating multiple matrix exponential derivatives (computational complexity ${\cal O}(K^{3})$) for each of the infinitesimal rates (${\cal O}(K^{2})$), leading to a total complexity of ${\cal O}(K^{5})$ operations and creating a severe computational bottleneck even for moderately-sized state spaces. To address this, we explore an approximation of the matrix exponential gradient that reduces the overall cost to only ${\cal O}(K^{2})$ (Didier et al., 2024), gaining a three orders-of-magnitude improvement. Crucially, using an approximate gradient does not hinder HMC convergence, as supported by the literature on HMC using surrogate trajectories (Li et al., 2019). Finally, we adapt the proposed methods to Bayesian phylogeographic inference following Magee et al. (2024), and demonstrate how the aforementioned approximation is even more critical for efficient sampling under flexible GP modeling assumptions.
@@ 2.1 Brief introduction to continuous-time Markov chains
More intuitively, if the CTMC $Y_{t}$ starts at some state $i$, it will leave such state after an exponentially distributed time with rate $q_{i}$ (time component) and will jump to a different state $j$ with probability $\frac{q_{ij}}{q_{i}}$ (location component). The exponential random variable describing the time spent in a state $i$ before leaving it has been assigned several names; in this work, we refer to it as the waiting time, $W_{i}$. When the CTMC is observed $N$ times, we denote the vector of observations by $\mathbf{y}=(y_{1},y_{2},\dots,y_{N})$. We refer to the times when each value $y_{i}$ is observed as $t_{i}$; we collect those values in the vector $\mathbf{t}=(t_{1},t_{2},\dots,t_{N})$.
@@ 2.2.1 Fully observed data
${p}\hskip-1.00006pt\left(\mathbf{y},\mathbf{t}\hskip-1.99997pt\mid\hskip-1.00006pt\mathbf{Q},\text{fully observed}\right)=\pi^{*}(y_{1})\prod_{i=2}^{N}\left[{\mathbb{P}}\hskip-1.99997pt\left(W_{y_{i-1}}=t_{i}-t_{i-1}\right)\frac{q_{y_{i-1},y_{i}}}{q_{y_{i-1}}}\right],$ (2)
@@ 2.2.2 Partially observed data
A more realistic scenario involves a partially observed CTMC. Specifically, the simplest case occurs when the state of the CTMC is known at time 0 and at time $t$, while its behavior throughout the interval $(0,t)$ remains unknown. This implies that an object of major importance is the finite-time transition probability of starting at a time 0 in state $i$ and being in state $j$ at a later time $t\geq 0$, that is:
@@
Collecting all such values in a matrix whose rows represent the starting states and columns the ending states, we obtain the finite-time transition probability matrix ${\mathbf{P}}\hskip-1.99997pt\left(t\right)=\left(P_{ij}(t)\right)_{ij}$. If we assume the identity matrix $\mathbf{I}$ as the initial condition of the system, that is, the system is in its current state without any transitions at time $t=0$, then it is easy to derive a closed-form formula for ${\mathbf{P}}\hskip-1.99997pt\left(t\right)$ based on the infinitesimal rate matrix $\mathbf{Q}$:
${\mathbf{P}}\hskip-1.99997pt\left(t\right)=e^{t\mathbf{Q}}:=\sum_{i=0}^{\infty}\frac{(t\mathbf{Q})^{i}}{i!},$ (4)
where $e^{t\mathbf{Q}}$ is the matrix exponential defined by its power series.
Now we consider specific cases where the data are only partially observed. We begin by assuming that we observe the data sequentially. Within this framework, we distinguish when the sampling times are known and when they are not. Then, we briefly introduce the more general case where observations are not assumed to be sequential.
@@
The simplest partially observed scenario arises when the sequence of observations includes both the states and the corresponding sampling times $(\mathbf{y},\mathbf{t})$, but there is no guarantee that all the jumps of the CTMC are observed. Then, we can write the joint distribution of $(\mathbf{y},\mathbf{t})$ using the Markov property and the finite transition probabilities:
${p}\hskip-1.00006pt\left(\mathbf{y},\mathbf{t}\hskip-1.99997pt\mid\hskip-1.00006pt\mathbf{Q}\right)=\pi^{*}(y_{1})\prod_{i=2}^{N}P_{y_{i-1},y_{i}}(t_{i}-t_{i-1}).$ (5)
With the exception of specific, mostly low-dimensional cases, such as the Jukes-Cantor (JK) substitution model (Jukes and Cantor, 1969) for DNA mutations, the values $P_{y_{i-1},y_{i}}(t_{i})$ are generally not available in closed-form. However, they can be computed by evaluating the matrix exponential for the specified time interval, which requires a computational effort of cubic-order (${\cal O}(K^{3})$).
@@
Suppose now the realizations in $\mathbf{y}$ are sequentially observed, but observation times are unknown. From Equation 4, it is clear that the waiting times and the infinitesimal rate matrix are jointly identifiable only up to a multiplicative factor. We resolve this by constraining $\mathbf{Q}$ and defining a time scale normalized by the expected number of jumps. Let $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{K})$ be a probability mass vector on $\mathbf{S}$ representing a priori state frequencies; $\boldsymbol{\pi}$ can represent $\mathbf{Q}$’s stationary distribution ($\boldsymbol{\pi}=\boldsymbol{\pi}^{*}$), but need not. Let $N_{\mathbf{Q}}(t)$ count the (random) number of jumps during $[0,t]$ with respect to $\boldsymbol{\pi}$. Its expected value is $\mathbb{E}\left[N_{\mathbf{Q}}(t)\right]=t\sum_{i\in\mathbf{S}}\pi_{i}q_{i}$, where $q_{i}=\sum_{j\neq i}q_{ij}$. To ensure this depends only on time, we rescale each entry of $\mathbf{Q}$ by the normalizing constant
@@
We then denote the normalized rate matrix $\frac{1}{\psi}\mathbf{Q}$ by $\mathbf{\Lambda}=(\lambda_{ij})_{ij}$. The CTMC defined by $\mathbf{\Lambda}$ has an expected number of jumps only dependent on time, $\mathbb{E}\left[N_{\mathbf{\Lambda}}(t)\right]=t$. $\mathbf{\Lambda}$ and $t$ become identifiable since any attempt to rescale the rate matrix globally is neutralized by the normalization constant. Here time is reinterpreted in terms of expected jumps: one time unit corresponds to the time necessary for one expected jump. In phylogenetics, this is called “evolutionary time” versus natural time. Assigning a prior ${p}\hskip-1.00006pt\left(\mathbf{t}\right)$ to the observation times, the joint distribution of the observations can be written as
@@
where ${p}\hskip-1.00006pt\left(\mathbf{y},\mathbf{t}\hskip-1.99997pt\mid\hskip-1.00006pt\mathbf{Q}\right)$ is computed as in Equation 5.
@@ 2.3 Prior on the infinitesimal rate matrix
The non-negativity constraints on the non-diagonal elements of the rate matrix $\mathbf{Q}$, if directly modeled, prevent the use of convenient prior distributions defined on the real line and complicate the use of gradient-based samplers. As is customary, we introduce a vector of real-valued auxiliary parameters $\boldsymbol{\theta}$ and link them to the rates by a non-negative differentiable one-to-one transformation $g(\cdot)$, such that:
@@
To relax the restrictive assumptions of the log-linear model employed in the previous literature (Pybus et al., 2012; Lemey et al., 2014; Zhao et al., 2016), we propose to model the relationship between the parameters and the covariate using a more flexible GP prior. This allows us to capture complex, nonlinear dependencies that a log-linear framework cannot adequately represent.
@@
The GP prior is a powerful Bayesian nonparametric prior that defines a distribution over more generic functions. Beyond accommodating nonlinearities, the GP enables a principled way to model uncertainty, as it provides not only point estimates but also credible intervals for inferred relationships. The GP is parameterized by a mean and a covariance function. The covariance function, also known as the GP kernel, defines the dependencies between function values at different input points, controlling both the smoothness and variability of the inferred function. By specifying an appropriate kernel, we can encode prior beliefs about the structure of the relationship, such as stationarity, periodicity, or varying degrees of smoothness. Common choices include the squared exponential kernel, which enforces smooth variations, and the Matérn kernel, which allows for more flexible roughness.
@@
If we consider a set of $P$ covariates $(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{P})$, then we can model the parameters $\theta_{ij}$ with an additive GP prior (Duvenaud et al., 2011), that is:
@@
where all $f_{p}$’s are independent and allowed to have different kernels $\mathbf{k}_{p}(\cdot,\cdot|\boldsymbol{\alpha}_{p})$. By exploiting the independence of the $f_{p}$’s, the expression can be integrated out analytically. In vector form, this implies that $\boldsymbol{\theta}$ follows a multivariate normal distribution,
$\boldsymbol{\theta}$ $\sim\mathcal{MVN}(\boldsymbol{0},\sum_{p=1}^{P}\mathbf{k}_{p}(\mathbf{x}_{p},\mathbf{x}_{p}^{\prime}|\boldsymbol{\alpha}_{p})).$ (11)
This framework can then be easily generalized to account for interactions between covariates (Duvenaud et al., 2011).